3) A student claims that two terms are enough to determine any sequence. For example 3,6, … means the sequence would be 3, 6, 9, 27, 12, 15, … How do you respond?
Answer by jim_thompson5910(35256) 
 The sequence could possibly be 3, 6, 9, 27, 12, 15, ... Notice how we add 3 to each term.
However, the sequence could also be 3, 6, 12, 24, 48, 96... Here we're multiplying each term by 2.
Or, you could multiply the each term by 3 and subtract 3. So 3*3-3=9-3=6
There are countless (possibly infinite) other ways to generate 3,6 ...
So there isn't enough information to determine the sequence given 3,6 ...
| |
|